# On the numerical stability of sketched GMRES

### v2

@@ 5 Adaptive restarting sGMRES
Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [BHMV2024, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
@@
Assume that Algorithm 1 is restarted, such that for restart number $i\geq 1$, we compute a basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ and a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (5) with high probability, where (7), (8), (11), (12), (13), and (27) are satisfied for $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$. Then there exists $k^{(i)}\leq n$ such that
$O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1.$ (45)
Assume that
$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1$ (46)
@@
$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert}.$
@@
Similar to the proof of [BHMV2024, Theorem 4.1], we will employ [BHMV2024, Lemma 4.2], which requires verifying the assumption [BHMV2024, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$. From Lemma 2, we have
@@
which gives [BHMV2024, (4.14)]. Together with [BHMV2024, Lemma 4.2], we can draw the conclusion. ∎
@@
Next, our aim is to propose an adaptive restarting strategy to recover from the instability of sGMRES caused by using a small $t$. The definition of $\Lambda_{1}^{(1)}$ in Theorem 2 relying on $\bar{\tau}_{k^{(i)}}$, which influences the reduced factor of the backward error in the $i$-th restart cycle, implies that care should be taken to ensure $\bar{\tau}_{k^{(i)}}$ does not get too large as the iteration progresses. In fact, as noted in [NT2024], due to the subspace embedding property (5), $\bar{\tau}_{i}$ can be inexpensively estimated as
@@
Algorithm 4 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive restarting
@@ 6 Numerical Experiments
In this section, we present numerical experiments that demonstrate the results of our stability analysis. All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM. We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [guttel2024sketch]. Our code is available at https://github.com/burkel8/stability_sGMRES. The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [balabanov2025randomized]).
We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-ssa, respectively, in the plots). Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [guttel2024sketch], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments. We also ran the experiments with sGMRES-ssa-greedy and obtained similar results, and thus do not include this method here. We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [guttel2024sketch] for more details on these methods. In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method. All comparisons to standard GMRES use GMRES with modified Gram-Schmidt. We do not use any preconditioner in the experiments, i.e., $M_{L}=M_{R}=I$. Finally, we drop the $\hat{.}$ notation in this section, as all quantities are computed quantities.
@@ 6.1 Illustration of the backward error bound
In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5. For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [NT2024]. We display the backward error, $\tau_{i}$ from (44), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration. These condition number approximations are denoted as $\kappa(B_{1:i})$ and $\kappa(AB_{1:i})$ in the plots.
@@ 6.1.3 Comparison between different sGMRES variants with and without restarting
In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [Davis2011] (also tested in [guttel2024sketch]). The right-hand side vector is randomly generated according to the standard normal distribution. We run $\texttt{nrestarts}=5$ cycles of restarted sGMRES with the truncated Arnoldi process, and sketch-and-select Arnoldi, with $m=150$ iterations per cycle, totaling $750$ iterations. We also run restarted GMRES with the same parameters. Additionally, we perform sGMRES with truncated Arnoldi and sketch-and-select Arnoldi with $m=750$ iterations (without restarts). All implementations of sGMRES use a truncation parameter $t=3$.
@@
6.2 sGMRES with the adaptive restarting strategy
In the final experiment, we test the adaptive restarting strategy discussed in Section 3. In particular, we test
@@
Figures 4 and 5 illustrate tests of three algorithms on matrices with different condition numbers $5.49\times 10^{3}$ and $9.64\times 10^{11}$.In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=10^{3}/m$.
@@
In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive restarting strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$. As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using an adaptive restarting strategy, tends towards the performance of restarted sGMRES.
Figure 5 clearly shows that, for $m=100$, our adaptive restarting strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$. This effect is even more pronounced for $m=400$. Conversely, when $m=50$, the backward error is limited by the small $m$, not $t$, since the behavior of restarted GMRES mirrors that of restarted sGMRES with a small $t$, which does not offer any improvement for this scenario. In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy performs similarly to standard GMRES. By comparing the experiments shown in Figures 4 and 5, we observe that restarted GMRES also requires a large $m$ in order to obtain a small backward error. According to Theorem 2, the reduced factor $\Lambda^{(i)}_{1}$ defined in (46) depends on $\bar{\tau}_{k^{(i)}}$, whose upper bound is determined by $\kappa(A)$. Therefore, for well-conditioned matrices such as fs_760_1, achieving $\Lambda^{(i)}_{1}\ll 1$ is easy, meaning that a relatively small $m$ typically suffices. In contrast, for ill-conditioned matrices, a significantly larger $m$ may be necessary.
Furthermore, in Figure 6, we combine the preconditioners generated by the MATLAB command ilu with the adaptive strategy to test sherman2, i.e., the same matrix tested in Figure 5. Due to the efficiency of the preconditioners, we can reduce the condition number of the preconditioned system and then employ much smaller $m$, that is, $m=5$, $10$. It can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
In general, sGMRES with our adaptive restarting approach often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
@@
Figure 4: Adaptive restarting experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
@@
Figure 5: Adaptive restarting experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
@@
Figure 6: Adaptive restarting experiment with preconditioning. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
@@ 7 Conclusions
In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis. However, in [guttel2024sketch, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned. To provide more insight into this finding from [guttel2024sketch], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$. We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive restarting strategy for sGMRES based on our analysis. Our numerical experiments indicate that the convergence of sketched GMRES with our adaptive restarting approach often performs comparably to standard GMRES and has the potential to result in substantial time savings.

### v3

@@ 5 Restarted sGMRES with adaptive truncation size
Furthermore, it can remove the influence of both $\kappa(M_{L})$ and $\kappa(Z_{1:j})$ on the backward stability of sGMRES, as shown in the following version of [3, Theorem 4.1] describing the stability of restarted sketched GMRES (Algorithm 2).
@@
Assume that Algorithm 1 is restarted every $m$ iterations with a randomized subspace embedding $S\in\mathbb{R}^{s\times n}$ satisfying (4) for the subspace $\mathbb{R}^{n\times m}$ with high probability. Also, assume that for each restart index $i\geq 1$ there exists $k^{(i)}\leq m$ such that the computed preconditioned basis $\hat{Z}_{1:k^{(i)}}^{(i)}\in\mathbb{R}^{n\times k^{(i)}}$ satisfies (7), (8), (11), (12), (13), (27) with $\varepsilon_{W}^{(i)}$, $\varepsilon_{sls}^{(i)}$, $\varepsilon_{S}^{(i)}$, $\varepsilon_{x}^{(i)}\leq O(\bm{u})$, and that
$\begin{split}&O(\bm{u})\sqrt{\frac{1-\epsilon}{1+\epsilon}}\kappa\bigl(\begin{bmatrix}\phi r^{(i)}&\hat{W}_{1:k^{(i)}}^{(i)}\end{bmatrix}\bigr)\geq 1,\qquad O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k^{(i)}}^{(i)})\leq\frac{1}{2},\\ &\text{and}\quad O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq\frac{1}{2}.\end{split}$ (45)
If
$\Lambda^{(i)}_{1}=\frac{1+\epsilon}{1-\epsilon}O(\bm{u})\kappa(M_{L})\bigl(1+\bar{\tau}_{k^{(i)}}\bigr)\ll 1,$ (46)
@@
$\bar{\tau}_{k^{(i)}}=\frac{\|\hat{Z}^{(i)}_{1:k^{(i)}}\rVert\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{y}^{(k^{(i)})}\rVert}{\|A\hat{d}^{(i)}\rVert},$
@@
Similar to the proof of [3, Theorem 4.1], we will employ [3, Lemma 4.2], which requires verifying the assumption [3, (4.14)], i.e., bounding $\|\hat{r}^{(i)}-A\hat{d}^{(i)}\rVert$ by $\|b-A\hat{x}^{(i)}\rVert$ and $\|b\rVert+\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}^{(i)}\rVert$. From Lemma 2, we have
@@
which gives [3, (4.14)]. Together with [3, Lemma 4.2], we can draw the conclusion. ∎
@@
$\frac{\|b-A\hat{x}\rVert}{\bigl\|A\bigr\rVert_{\mathsf{F}}\|\hat{x}\rVert+\|b\rVert}\leq O(\bm{u}).$
@@
Next, our aim is to propose an adaptive strategy to recover from the instability of sGMRES caused by using a small $t$. The definition of $\Lambda_{1}^{(1)}$ in Theorem 2 relying on $\bar{\tau}_{k^{(i)}}$, which influences the reduced factor of the backward error in the $i$-th restart cycle, implies that care should be taken to ensure $\bar{\tau}_{k^{(i)}}$ does not get too large as the iteration progresses. In fact, as noted in [10], due to the subspace embedding property (4), $\bar{\tau}_{i}$ can be inexpensively estimated as
@@
Algorithm 4 One cycle of preconditioned sketched GMRES (sGMRES) with adaptive truncation size
@@ 6 Numerical Experiments
In this section, we present numerical experiments that demonstrate the results of our stability analysis. All experiments were performed in MATLAB R2023a on a Windows 11 HP laptop with an 11th Gen Intel(R) Core processor with 2.80 GHz and 8 GB of RAM. We used (with modifications) the open source code provided by Güttel and Simunec as part of their recent work [6]. Our code is available at https://github.com/burkel8/stability_sGMRES. The sketching operator used in all experiments is the subsampled random Hadamard transform (see, e.g.,  [1]).
We consider solving a linear system $Ax=b$ with sGMRES using both the truncated Arnoldi process and the sketch-and-select Arnoldi method (sGMRES-trunc and sGMRES-ssa, respectively, in the plots). Although a few variants of the sketch-and-select Arnoldi algorithm have been proposed in [6], we use the sGMRES-ssa-pinv method, as it was shown to produce the best conditioned basis in their experiments. We also ran the experiments with sGMRES-ssa-greedy and obtained similar results, and thus do not include this method here. We note that the different variations of sketch-and-select Arnoldi differ in how the $t$ vectors with which to orthogonalize against are chosen, and we refer to [6] for more details on these methods. In these experiments, we always choose the same value of $t$ in the sketch-and-select Arnoldi algorithm that we use as the truncation parameter in the truncated Arnoldi method. All implementations of standard GMRES use the modified Gram-Schmidt Arnoldi process. We do not use any preconditioner, i.e., $M_{L}=M_{R}=I$, in the experiments except for Figure 6. Finally, we drop the $\hat{\cdot}$ notation in this section, as all quantities are computed quantities.
@@ 6.1 Illustration of the backward error bound
In this subsection, we use three different examples to illustrate the results of our analysis shown in Sections 4 and 5. For all sketching techniques we take $s=2(m+1)$, where $m$ is the maximum length of each Arnoldi cycle, as suggested in [10]. We display the backward error, $\tau_{i}$ from (44), and the condition numbers of $B_{1:i}$ and $AB_{1:i}$ (approximated using the condition numbers of $SB_{1:i}$ and $SAB_{1:i}$, respectively) at each iteration. These condition number approximations are denoted as $\kappa(B_{1:i})$ and $\kappa(AB_{1:i})$ in the plots.
@@ 6.1.3 Comparison between different sGMRES variants with and without restarting
In the third example, the matrix $A$ is the Norris/stomach matrix of size $n=213,360$ obtained from the SuiteSparse matrix collection [4] (also tested in [6]). The right-hand side vector is randomly generated according to the standard normal distribution. We run $\texttt{nrestarts}=5$ cycles of restarted sGMRES with the truncated Arnoldi process, and sketch-and-select Arnoldi, with $m=150$ iterations per cycle, totaling $750$ iterations. We also run restarted GMRES with the same parameters. Additionally, we perform sGMRES with truncated Arnoldi and sketch-and-select Arnoldi with $m=750$ iterations (without restarts). All implementations of sGMRES use a truncation parameter $t=3$.
@@
6.2 sGMRES with adaptive truncation size
In the final experiment, we test the adaptive strategy discussed in Section 3. In particular, we test
@@
  • restarted GMRES with $m=t_{\max}$: restart every $t_{\max}$ iterations, where $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
@@
Figures 4 and 5 illustrate tests of these algorithms on matrices fs_760_1 and sherman2, whose condition numbers are $5.49\times 10^{3}$ and $9.64\times 10^{11}$, respectively. In each example, the right-hand side is generated randomly following the standard normal distribution, with $m$ values of $50$ and $100$, and $\texttt{nrestarts}=1.5\cdot 10^{3}/m$. It should be noted that restarted GMRES with $m=t_{\max}$ is examined primarily to enable a direct comparison between restarted sGMRES with adaptive $t$ and restarted GMRES with similarly low orthogonalization costs. In Figure 5, we omit the result for “restarted GMRES with $m=t_{\max}$” because it coincides with that of “restarted GMRES”.
@@
In Figure 4, both restarted GMRES and sGMRES achieve small backward errors, with restarted sGMRES using the adaptive strategy accelerating convergence compared to a fixed $t$ approach by using a slightly larger $t$. As shown in the middle subfigure of Figure 4, it can be expected that the performance of restarted sGMRES, using adaptive truncation size, tends towards the performance of restarted sGMRES. To further illustrate the potential performance advantage of our adaptive strategy, we also compare the number of floating-point operations required for orthogonalization in Table 1. From the table, we observe that restarted sGMRES, both with and without the adaptive strategy, entails a comparable but substantially lower orthogonalization cost than restarted GMRES.
@@
Figure 4: Adaptive truncation size experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for fs_760_1, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3 and $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
We next present a difficult example for restarted sGMRES in Figure 5. In this situation, restarted sGMRES struggles to attain a small backward error. As indicated in Theorem 2, the attainable backward error is affected by two conditions, namely (45) and (46). For $m=50$ in Figure 5, the value of $m$ is too small for (45) to be satisfied. Consequently, even if we use restarted sGMRES with the adaptive strategy or restarted GMRES, both of which ensure a small $\tau$ so that (46) holds, restarting still fails to improve the attainable backward error. In contrast, when $m=100$ and $m=400$, the larger values of $m$ allow (45) to be satisfied. In this situation, $\tau$, associated with (46), becomes the dominant factor determining the attainable backward error. Figure 5 clearly shows that, for $m=100$, our adaptive strategy can enhance the poor backward error that arises from a small $t$ by adaptively increasing $t$ to reduce $\tau$ such that (46) is satisfied. This effect is even more pronounced for $m=400$. In this case, $t$ is repeatedly doubled up to $m$ during the early stage, thus restarted sGMRES with the adaptive strategy utilizes a comparable orthogonalization cost as standard GMRES. By comparing the experiments shown in Figures 4 and 5, we observe that restarted GMRES also requires a large $m$ in order to obtain a small backward error. According to Theorem 2, $O(\bm{u})\sqrt{\frac{1+\epsilon}{1-\epsilon}}\kappa(\hat{W}_{1:k^{(i)}}^{(i)})\leq 1/2$ in (45) depends on $\kappa(\hat{W}_{1:k^{(i)}}^{(i)})$, whose upper bound is determined by $\kappa(A)$ and $\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})$. This implies that, in the worst case, it may be necessary to require $O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq 1/(2\kappa(A))$ in order to guarantee (45). Consequently, for well-conditioned matrices such as fs_760_1, $\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})$ can be relatively large while still satisfying $O(\bm{u})\kappa(\hat{Z}_{1:k^{(i)}}^{(i)})\leq 1/(2\kappa(A))$, meaning that a relatively small truncation size usually suffices. In contrast, for ill-conditioned matrices, a significantly larger small truncation size may be required.
@@
Figure 5: Adaptive truncation size experiment. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3.
Furthermore, in Figure 6, we combine the preconditioners generated by the MATLAB command ilu with the adaptive strategy to test sherman2, i.e., the same matrix tested in Figure 5. Due to the efficiency of the preconditioners, a small $m$, that is, $m=50$, is sufficient to achieve convergence. From Figure 6 together with Table 1, it can be seen that when preconditioning is used, restarted sGMRES with and without the adaptive strategy have similar behavior and require fewer orthogonalizations than standard GMRES.
In general, sGMRES with adaptive truncation size often performs comparably to standard GMRES regarding the accuracy, and has the potential to result in substantial time savings especially for the well-conditioned case.
@@
Figure 6: Adaptive truncation size experiment with preconditioning. From left to right, we plot the backward error, the value of $\bar{t}$, and the value of $\tau$ for sherman2, where $\bar{t}=\max(t,i)$ represents the number of columns against which $B_{i}$ needs to perform orthogonalization in Line 2 of Algorithm 3 and $t_{\max}$ denotes the maximum truncation size employed in the “restarted sGMRES with adapt tt” method.
Table 1: Comparison of the relative number of floating-point operations used in orthogonalization. In the table, we report the ratios of the number of floating-point operations required for the orthogonalization of “restarted sGMRES,” “restarted sGMRES with adapt tt,” and “restarted GMRES with $m=t_{\max}$” relative to that of “restarted GMRES,” respectively.
m
restarted
GMRES
restarted
sGMRES
restarted
sGMRES
with adapt $t$
restarted
GMRES
with $m=t_{\max}$
fs_760_1 without pre 50 1 0.3072 0.5959 1.7979
100 1 0.5468 0.4485 3.4323
fs_760_1
with pre
50 1 0.1250 0.1250 0.1883
@@ 7 Conclusions
In this paper, we performed a backward stability analysis of sketched GMRES and showed how the backward error is largely influenced by the condition number of the generated Krylov basis. However, in [6, Figure 9], it was demonstrated that the backward error of sGMRES can still decrease even when the basis becomes ill-conditioned. To provide more insight into this finding from [6], we proposed a sharper bound that relies on $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ rather than $\kappa(\hat{B}_{1:k})$. Nevertheless, it remains unclear why $\|\hat{B}_{1:k}\rVert\|\hat{y}^{(k)}\rVert/\|\hat{x}^{(k)}\rVert$ itself can even decrease—leading to a reduction in the backward error of sGMRES—when the basis becomes ill-conditioned. We then showed how restarting can be used to stabilize a sketched GMRES implementation, and proposed an adaptive strategy to choose proper truncation sizes for sGMRES based on our analysis. Our numerical experiments indicate that the convergence of sketched GMRES with adaptive truncation size often performs comparably to standard GMRES and has the potential to result in substantial time savings.